OpenStudy (anonymous):

joe co. in 2006 had sales of $800,000. in 2007 sales were up 75 percent . calculate the sales for 2007 ?????

OpenStudy (anonymous):

800,000=100% 1%= 800,000/100 = 8,000 75% = 8,000 x 75 =600,000 therefore sales for 2007 = 800,000 + 600,000=1,400,000 presuming of course that sales were up 75% from 2006!